Calculate the molality of 6.5 M aqueous solution of an acid ,HA with a density of 0.888 g cm^-3.Given the molar mass of acid is 98.0 g mol^-1


1
Expert's answer
2020-09-06T15:00:02-0400

First we need to find the mass of this solution.

We know that "Density(p) = Mass(m)\/Volume(V)"

Thus making m the subject of the formula gives

"m=pV"

(p=0.888g/cm-3 and V=1L=1000cm3)

Thus, "m=(0.888g\/cm^3) x (1000cm^3) = 888g" (mass of the entire solution)


Secondly, we need to get the mass of water in the solution.

We have 6.5M HA, what is the mass of HA are there?

Therefore, "1 mol of HA=98.0g"


"6.5 mol of HA =?"

By cross multiplication we get;

"\frac{6.5 mol x 98.0g}{1} = 637g" (mass of HA in the solution)


"\therefore mass of waterin the solution = 888g - 637g = 251g =0.251kg"

Calculate molality using the formula


"Molality=\frac{moles of solute}{mass in Kg of solvent}"


"= \frac{6.5 mol}{0.251Kg}"


"= 25.896 mol\/Kg"